OpenStudy (anonymous):

Find the slope-intercept form of the equation of the line that passes through the given point and has the indicated slope. Point Slope (−1, 2) m = −2

OpenStudy (hoblos):

y-2 = -2(x+1) y=-2x

OpenStudy (anonymous):

point slope first, then solve for y to get slope intercept formula